OpenStudy (anonymous):

Solve the system by substitution. 2x-y+z=-4 z=5 -2x+3y-z=-10

OpenStudy (anonymous):

x=-8 y=-7

OpenStudy (campbell_st):

put z = 5 into equations 1 and 3 2x - y + 5 = -4 or 2x - y = -9 equation 1 -2x + 3y - 5 = -10 -2x + 3y = - 5 equation 2 rewrite equation 1 2x = y - 9 substitute it into equation 2 -(y - 9) + 3y = -5 solve for y...